SPHERICAL TRIGONOMETRY

  1. solve for the three vertices of a spherical triangle ABC if the sides measures a=68degree48', b=53degree15' and c=46degree30'.
  2. solve for these right spherical triangle with parts side c=69degree25' and angle B=63degree25' note that this is a right spherical triangle. (HINT: use the Napier's circle to determine what can be solved first by the given parameters.)
  3. solve for the area of the spherical triangle in problem 1 and 2 if the sphere has a radius of 30m.
  4. *show solution*
1
Expert's answer
2022-01-17T18:34:57-0500

compilation of the THREE questions


(1) we shall solve the first question using the Haversine formula

"havA=\frac{hav \hspace{0.1cm} a - hav(b-c)}{sin b* sin c}"


We convert the give parameters to degrees

"a=68°41^{\prime}=68.8°\\b=54°15^{\prime}=53.25°\\c=46°30^{\prime}=46.5°"


We know that

"hav(\theta)=sin^{2}(\frac{\theta}{2})"


We shall then find for the parameters a, b and c

"hav(a)=hav(68.8)=sin^{2}(\frac{68.8}{2})\\\hspace{1.cm}=sin^{2}(34.5)=0.5649^{2}=0.3192"


"hav(b)=hav(53.25)=sin^{2}(\frac{53.25}{2})\\\hspace{1.cm}=sin^{2}(26.625)=0.4481^{2}\\\hspace{1cm}=0.2008"


"hav(c)=hav(46.5)=sin^{2}(\frac{46.5}{2})\\\hspace{1.cm}=sin^{2}(23.25)=0.3947^{2}\\\hspace{1cm}=0.1558"



Also, we have

"hav(b-c)=hav(53.25-46.5)\\\hspace{1cm}=hav(6.75)=sin^{2}(\frac{6.75}{2})\\\hspace{1.cm}=sin^{2}(3.375)=0.589^{2}\\\hspace{1cm}=0.00347"


"hav(a-c)=hav(68. 8-46.5)\\\hspace{1cm}=hav(22.3)=sin^{2}(\frac{22.3}{2})\\\hspace{1.cm}=sin^{2}(11.15)=0.1934^{2}\\\hspace{1cm}=0.0374"


"hav(a-b)=hav(68.8-53.25)\\\hspace{1cm}=hav(15.55)=sin^{2}(\frac{15.55}{2})\\\hspace{1.cm}=sin^{2}(7.775)=0.1353^{2}\\\hspace{1cm}=0.0183"



The sin of the parameters a, b and c

"sin(a)=sin(68.8°)=0.9323\\sin(b)=sin(53.25°)=0.8013\\sin(c)=sin(46.5°)=0.7254"


Now

Using the Haversine formula to solve for havA, havB and havC

"havA=\frac{0.3192-0.00347}{0.8013*0.7254}=0.5432\\"



"havB=\frac{0.2008-0.00347}{0.9323*0.7254}=0.2416"



"havC=\frac{0.1558-0.0183}{0.9323*0.8013}=0.1840"



"havA=0.5432\\\implies0.5432=sin^{2}(\frac{\theta}{2})\\\implies\sqrt{0.5432}=sin(\frac{\theta}{2})\\\implies0.7370=sin(\frac{\theta}{2})\\\implies sin^{-1}(0.7370)=\frac{\theta}{2}\\\implies 47.4764=\frac{\theta}{2}\\\implies \theta=94.95°
\\A=94.95°"



"havB=0.2416\\\implies0.2416=sin^{2}(\frac{\theta}{2})\\\implies\sqrt{0.2416}=sin(\frac{\theta}{2})\\\implies0.4915=sin(\frac{\theta}{2})\\\implies sin^{-1}(0.4915)=\frac{\theta}{2}\\\implies 29.44=\frac{\theta}{2}\\\implies \theta=58.88°
\\B=58.88°"




"havC=0.1840\\\implies0.1840=sin^{2}(\frac{\theta}{2})\\\implies\sqrt{0.1840}=sin(\frac{\theta}{2})\\\implies0.4289=sin(\frac{\theta}{2})\\\implies sin^{-1}(0.4289)=\frac{\theta}{2}\\\implies 25.4011=\frac{\theta}{2}\\\implies \theta=50.80°
\\C=50.80°"




(2) We have a right spherical triangle with

"c=69\u00b025`, B=63\u00b025`, A=90\u00b0"


In the first

"Sin(c)=tan(b) * tan(90-B)"

"\therefore sin(c)=tan(b) * cot(B)\\\therefore tan(b)=\frac{sin(c)}{cot(B)}=sin(c) * tan(B)"

"=sin(69.42)-tan(63.42)\\=1.8711"

"\therefore tan(b)=1.8711\\b=tan^{-1}1.8711\\b=61.88"



In the second

"sin(90-C)=cos(c) * cos(90-B)\\\therefore cos(C)= cos(c) * sin(B)\\\therefore cos(C)=cos(69.42) * sin(63.42)"

"=0.3143"

"C=cos^{-1}0.3143\\C=71.68°"



In the third

"sin(90-B)=tan(c) * tan(90-a)\\\therefore cos(B)=tan(c) * cot(a)\\\therefore cos(B)=\frac{tan(c)}{tan(a)}\\\therefore tan(a)=\frac{tan(c)}{tan(B)}"

"=\frac{tan(69.42)}{cos(63.42)}=\frac{2.6633}{0.4474}=5.9528"


"a=tan^{-1}(5.9528)\\a=80.46°"



(3) we shall calculate the Area of the spherical triangle in (1) and (2)

To calculate the Area for (1)

Using the formula

"A=\frac{\pi R^{2} E}{180}"

Where E is the excess and we have

"E= A+B+C - 180\u00b0"

To calculate the Excess for question (1)

"E=94.95°+58.88°+50.80° - 180°\\E=24.63"


Now

"A=\frac{(3.142) (30)^{2} (24.63)}{180}"


"A=\frac{(3.142) (900) (24.63)}{180}"


"A=\frac{69648.714}{180}\\A=386.9"


Also, for question (2)

"E=90°+63.42°+71.68°-180°\\E=45.1"


"A=\frac{\pi R^{2} E}{180}"


"A=\frac{(3.142) (30)^{2} (45.1)}{180}"


"A=\frac{(3.142) (900) (45.1)}{180}"


"A=\frac{127533.78}{180}\\A=708.521"